The problem states, the Museum charges 21 for adults, $14 for seniors, and $5 for children under 12 for tours. This past sunday, the paid revenue was $9417 for 735 tickets sold. There were 43 more children than adults. How many children attended?
Answer by stanbon(75887)

The problem states, the Museum charges 21 for adults, $14 for seniors, and $5 for children under 12 for tours. This past sunday, the paid revenue was $9417 for 735 tickets sold. There were 43 more children than adults. How many children attended?
---------------------------
Quantity Equation: a + s + c = 735
Value Equation: 21a +14s + 5c = 9417
Comparison Eq: c = a + 43
----------------------------------------
Rearrange:
a + s + c = 735
21a +14s + 5c = 9417
a + 0 - c = -43
-----------------------
solve this system by any method you know to get:
a = 243 (# of adults)
s = 206 (# of seniors)
c = 286 (# of children)
=============================
Cheers,
Stan H.